A rectangular patio is 12 feet long and 8 feet wide. The owner wants to increase both dimensions by the same amount resulting in a patio that has an area of 221 sq. ft. Write and solve an equation to find out how much the owner should increase each dimension.
Answer by mananth(16946)

width =8
length = 12
increase both by same amount x
8 +x & 12 +x
Area = L*W
Area = 221 sq. ft
221= (8+x)(12+x)
221 = 96 +8x+12x+x^2
221-96=x^2+20x
125=x^2+20x
x^2+20x-125=0
x^2+25x-5x-125=0
x(x+25)-5(x+25)=0
(x+25)(x-5)=0
x=5 OR -25
neglect the negative
x= 5 the width increase
m.ananth@hotmail.ca